OpenStudy (anonymous):

Convert 65° into radians

OpenStudy (ash2326):

Do you know that \[\pi \ radians=180\ degrees\]?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

so \[3.14\ radians=180\ degrees\] \[\frac{3.14}{180}\ radians=1\ degree\] so \[1\ degree=\frac{3.14}{180}\ radians\] hence \[65\ degrees =65\times \frac{3.14}{180}\ radians\]

OpenStudy (ash2326):

Can you do the rest?

OpenStudy (anonymous):

oh!! okay now i get it. thanks:)

OpenStudy (ash2326):

welcome:D